Ari dropped a golf ball from her second story window. The ball starts from rest

and hits the sidewalk 1.5 s later with a velocity of 14.7 m/s. Find the average

acceleration of the golf ball.



1
Expert's answer
2021-10-05T15:43:51-0400
"a_{avg}=\dfrac{v-v_0}{t}=\dfrac{14.7\ \dfrac{m}{s}-0}{1.5\ s}=9.8\ \dfrac{m}{s^2}."